OpenStudy (anonymous):

If A= 2i+4j+5k and B=-2i+2j+k. then A.B is ??

OpenStudy (anonymous):

A.B=(2i+4j+5k).(-2I+2J+K) \[=\left( 2 \right)\left( -2 \right)+\left( 4 \right)\left( 2 \right)+\left( 5 \right)\left( 1 \right)\] =-4+8+5 =9